OpenStudy (anonymous):

Which is the most reasonable estimate for seventy-five percent of one thousand one hundred eighty-two? nine hundred nine hundred fifty one thousand one thousand fifty

OpenStudy (anonymous):

take 75% of 1200 which is 3/4 of 1200 = 900

OpenStudy (anonymous):

75% = 3/4

jigglypuff314 (jigglypuff314):

1182 * 0.75 =

OpenStudy (anonymous):

thx :D :) :* 6-6 ^_^ @_@ !_! @wolfgirl14

OpenStudy (anonymous):

i mean @jigglypuff314

jigglypuff314 (jigglypuff314):

lol your welcome :)

OpenStudy (anonymous):

:A :B :C :D :E :F :G :H :I :J :K :L: :M :N :O :P :Q :R :S :T :U :V :W :X :Y :Z @jigglypuff314